OpenStudy (anonymous):

if f(x)=6x+3 and g(x)=5x+4 what is f(g(x))? A.30x-21 B.30x-12 C.30x+1 D.30x+11

OpenStudy (anonymous):

Plug g(x) into x. f(g(x))=6*g(x)+3=6(5x+4)+3

OpenStudy (raden):

no options be right answer

OpenStudy (raden):

except if f(x) = 6x - 3

OpenStudy (anonymous):

I'm guessing g(x) is actually 5x-4.

OpenStudy (anonymous):

@Anonymous1921 it is, I wrote it wrong D:

OpenStudy (raden):

ah, f(x) = 6x - 3, cant be right answer too --" yeah, the g(x) = 5x-4.